Im really confused since our teacher doesnt explain well, although thats not really a good reason.
(√4:(7x^2y))/(√3:(4x^5y^2))
can you show me the answer and solution if possible? thanks!

Answer by stanbon(75887)

(√4:(7x^2y))/(√3:(4x^5y^2))
--
= 2/(7x^2y) / sqrt(3)/(4x^5y^2)
----
Invert the denominator and multiply to get:
=====
= 2/(7x^2y) * (4x^5y^2)/sqrt(3)
-----
= (2/sqrt(3))[28x^7y^3]
-----
= (2/3)sqrt(3)(28x^7y^3)
===============================
Cheers,
Stan H.
==============